Grade 12Physical Chemistry

105ml of pure water(4oC)is saturated with NH3 gas,producing a solution of density 0.9g/ml.if this solution contains 30% of NH3 by weight , calculate its volume

nikesh
10 Years agoGrade 12

1 Answer

K10
9 Years ago
Let V (mL) be the unknown volume of density 0.9 g/mL. Density = mass / volume. Therefore, mass of solution = 0.9V g. Given 105 mL of water of density 1 g/mL, its mass = 1 (g/mL) * 105 (mL) = 105 g. But mass of water is also equal to 70% of mass of solution, i.e. (70/100) * 0.9V g. Equating masses of water : 105 = (70/100) * 0.9V Therefore, V = 105 * 100 / (0.9 * 70) = 166.7 mL.